OpenStudy (anonymous):

Simplify (1-hk^-1)/(h^-1-k^-1)

OpenStudy (anonymous):

\[\frac{ 1-hk^{-1} }{ h ^{-1}-k ^{-1} }\]

OpenStudy (anonymous):

\[\frac{ 1- \frac{ h }{ k } }{ \frac{ 1 }{ h }-\frac{ 1 }{ k } }\] Try looking at it like this and see if you can solve.

OpenStudy (anonymous):

okay, thank you!